OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -9 - 3 + 3 + 9 + ... + 81

OpenStudy (anonymous):

what is the pattern you see?

OpenStudy (anonymous):

I can't seem to see any.

OpenStudy (anonymous):

adding 6

OpenStudy (anonymous):

:O now I see it lol. I was thinking harder than I needed to lol

OpenStudy (anonymous):

so the answer would be

OpenStudy (anonymous):

instead of that 15 would it be infinity?

OpenStudy (anonymous):

heheh summations are meant to look difficult but i've always found them fun and easy!

OpenStudy (anonymous):

well, you are stopping at 81?

OpenStudy (anonymous):

oh so it would be 15.

OpenStudy (anonymous):

very well set up! Just in case you are not aware, summations can have MANY different answers so... often you will see your way is not the same as the back of the book. Doesn't mean you are wrong though! I remember WAY more often than not, my approach wasn't the same

OpenStudy (anonymous):

Do you mind helping me with one more I am confused with?

OpenStudy (anonymous):

sure thing

OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. 64 + 81 + 100 + 121 + ... + n2 + ...

OpenStudy (anonymous):

see a pattern by any chance?

OpenStudy (anonymous):

sorta. It's the difference of the two numbers + 2

OpenStudy (anonymous):

oh? can you explain?

OpenStudy (anonymous):

81-64 = 17 100-81=19 121-100=21 ...

OpenStudy (anonymous):

aha! that looks right! now we need an equation that works. i like to pick what i want my lower bound to be first and make an equation that will work with it

OpenStudy (anonymous):

the n is adding 2 to itself right?

OpenStudy (anonymous):

(n+2) looks like a pretty good candidate to have in your equation.

OpenStudy (anonymous):

oooh that also!

OpenStudy (anonymous):

awesome! thanks